2. Minimize: C-2x+5) Subject to: 4x+y 240 2? + ? 230 x+3y 2 30 x20,y20 ) Type in the corner points found and their corresponding Cost b) What is the minimum cost?
Answer #1

minimize

C=2x + 5y

subjected to

4x+y\ge 40\:

\:2x+y\ge 30\:

+3y2 30

plot on graph

40 (0,40) 20 (5,20) (12,6) (30,0) 20

corner points are

(30,0)......2x+5y......=60

(12,6)......2x+5y......=54.......minimum

(5,20)......2x+5y......=110

(0,40)......2x+5y......=200

at point (12,6) function is minimum

minimum cost is 54